OpenStudy (anonymous):

how exactly do you know what the level curves of a function are? for example: f(x,y)=x^2+y

OpenStudy (anonymous):

Equate the function to a constant. This equation will describe your curves.

OpenStudy (anonymous):

So in your example: x^2+y=c y=c - x^2

OpenStudy (anonymous):

so once you set the equation to a constant,"c', you solve for y? plus where can i find like a list of the equations for like a hyperbola and ellipse? i set the equation to a constant but i get stuck when it comes to recognizing the graph

OpenStudy (anonymous):

You don't have to solve for y, if your function is x^2+y^2, the curves will described by x^2+y^2=c. You should recognize this being a circle.

OpenStudy (anonymous):

here's a small list: http://en.wikipedia.org/wiki/Conic_section#Features